OpenStudy (anonymous):

Can someone help me. I need help. Write a polynomial function of the nth degree that has the given real or complex roots. n=3; x= -2, x=3i

OpenStudy (anonymous):

Please I will give a medal!!!

OpenStudy (anonymous):

help me please

OpenStudy (fibonaccichick666):

so first what can you write down for a zero at x=2?

OpenStudy (anonymous):

The zero is 2 I think I got it is this correct? x^3 - x^2-6x-3ix^2+3ix + 18i

OpenStudy (fibonaccichick666):

uh... you should leave it un-muiltiplied out

OpenStudy (campbell_st):

an easy way to look at this problem is start with the complex roots. you have \[x = \pm 3i\] if you square both sides of the equation you get \[x^2 = 9i^2\] and you should know \[i^2 = -1\] so then \[x^2 = -9\] which can be written as a quadratic factor of \[(x^2 + 9) = 0\] you linear factor comes from x = -2 is a zero so x + 2 is a factor combining the factors gives \[(x + 2)(x^2 + 9) = 0\] your task is to now expand to get the cubic